If 10g of phosphine is made to react with 15g of chlorine gas , calculate the percentage yield if 5.5g of phosphorous pentachloride was experimentally produced
1
Expert's answer
2019-04-02T05:58:53-0400

2 PH3 + 8 CL2 = 2 PCL5 + 6 HCL

M(Cl2) = 71 g/mol

M(PCL5) = 208.24 g/mol

M(PH3) = 40.0 g/mol

For equation we can see, that PH3 is an excess. So the calculations are carried out on chlorine.

568 g (CL2) = 416.48 g (PCL5)

15 g (Cl2) = x g (PCL5)

x = 416.48*15/568 = 11.0g (PCL5)

For experiment yield is 5.5g

Percentage yield = 5.5/11.0*100% = 50%

Answer: 50%